OpenStudy (anonymous):

Could use some help with this -5+6(7-5r)=-r-21

OpenStudy (anonymous):

r=2

OpenStudy (agreene):

-5+6(7-5r)=-r-21 -5+42-30r=-r-21 37-30r=-r-21 58=29r 2=r

OpenStudy (anonymous):

Thanks in advance for helping me out in this one.